Calculate the amount of 
1) H2SO4 to prepare 500cm^3 of M/50 H2SO4 solution
2) Potash alum to prepare 250ml of M/20 potash alum solution.

Grade:11

3 Answers

Khimraj
5 years ago
​i) let mass of H2SO4 is x then
mole/volume(L) = M/50
(x/98)/0.5) = 1/50
x = 49/50 gm = 0.98gm
Hope it clears.
Khimraj
5 years ago
 
i​i) let mass of potash alum is x then
mole/volume(L) = M/20
(x/949)/0.24) = 1/20
So x = 11.39gm
Hope it clears.
Khimraj
5 years ago
i​i) let mass of potash alum is x then
mole/volume(L) = M/20
(x/949)/0.25) = 1/20
So x = 11.86gm
Hope it clears.